# TriG-NER: Triplet-Grid Framework for Discontinuous Named Entity Recognition

### v2

@@ 6. Conclusion
Appendix A Comprehensive Metric Scores
We provide the F1, Precision, and Recall scores from our overall best-performing model in Table A1. In Table A2, we present the performance scores from the model setup that scores highest for the discontinuous entities only (DiscEnt). We observe significantly higher scores for discontinuous entities for the best DiscEnt model with 1.66%, 2.92%, and 4.83% for CADEC, ShARe13, and ShARe14, respectively. However, despite not having the best overall scores in our experiments, the best DiscEnt models still outperform all of the baselines for the CADEC and ShARe14 datasets and are comparable to our overall best model highlighting the ability of our framework to extract discontinuous entities through word-pair triplets.
Table A1. Complete performance scores from the best-performing overall model for sentences with at least one discontinuous entity (DiscSent) and for discontinuous entities only (DiscEnt).
@@ Appendix A Comprehensive Metric Scores
Table A2. Complete performance scores from the best-performing discontinuous entity model for the overall dataset, for sentences with at least one discontinuous entity (DiscSent), and for discontinuous entities only (DiscEnt).
@@
Appendix B Token Gap Analysis
Figure B1 shows the difference in token gaps between CADEC, ShARe13, and ShARe14. CADEC generally shows shorter gaps between spans for discontinuous entities, while the ShARe datasets have wider gaps despite having shorter entities. These differences present unique challenges for extracting discontinuous entities in each dataset, highlighting the need for a flexible and adaptable solution like our proposed framework.
Figure B1. Distribution of token gaps of discontinuous entities.
Appendix C Hyperparameter Study
We investigate the effect of different hyperparameter values on our best overall model. In Table C1, we test different learning rate values for the Adam optimiser and find that the optimal learning rate value for our framework is 5e-04.
Table C1. Comparison of learning rates. Bold indicates best scores while underline shows next best.
@@ Appendix C Hyperparameter Study
Appendix D Best-found Parameter Setup
Table D1. Parameter setup for the best model based on overall performance scores for each dataset.
@@ Appendix D Best-found Parameter Setup
Appendix E Large Language Model Prompts
Table E1. Prompt templates used for large language models.
@@ Appendix E Large Language Model Prompts
[…]
Table E2. Variables and examples used for prompt engineering for each dataset.
@@
Appendix F Case Studies
Figure F1. Case study for CADEC comparing results from trained models using our framework and a baseline and from zero and few-shot CoT prompt engineering using LLMs. The sample prompt provided follows the few-shot CoT template. All prompt templates are provided in Appendix E.
Figure F2. Case study for CADEC comparing results from trained models using our framework and a baseline and from zero and few-shot CoT prompt engineering using LLMs. The sample prompt provided follows the few-shot CoT template. All prompt templates are provided in Appendix E.
Figure F3. Case study for ShARe13 comparing results from trained models using our framework and a baseline and from zero and few-shot CoT prompt engineering using LLMs. The sample prompt provided follows the few-shot CoT template. All prompt templates are provided in Appendix E.
Figure F4. Case study for ShARe14 comparing results from trained models using our framework and a baseline and from zero and few-shot CoT prompt engineering using LLMs. The sample prompt provided follows the few-shot CoT template. All prompt templates are provided in Appendix E.

### v3

@@ 6. Conclusion
Acknowledgements.
This study was supported by funding from the Google Award for Inclusion Research Program (G222897).
Appendix A Dataset Statistics
We provide a breakdown of relevant statistics for each dataset on Table A1. CADEC is the smallest dataset with 7,597 sentences while ShARe14 is the largest with 34,618 sentences. All datasets have only around 10% of discontinuous entities.
Table A1. Data statistics
CADEC ShARe13 ShARe14
Total Sentences 7,597 18,767 34,618
Total Entities 6,318 11,148 19,073
Continuous Entities 5,639 10,060 17,417
- Percentage 89.25% 90.24% 91.32%
- Number of tokens 1-36 1-9 1-9
Disc. Entities 679 1,088 1,658
- Percentage 10.75% 9.76% 8.68%
- Number of tokens 2-13 2-7 2-7
- Start-End Distance 3-20 3-23 3-23
Appendix B Comprehensive Metric Scores
We provide the F1, Precision, and Recall scores from our overall best-performing model in Table B1. In Table B2, we present the performance scores from the model setup that scores highest for the discontinuous entities only (DiscEnt). We observe significantly higher scores for discontinuous entities for the best DiscEnt model with 1.66%, 2.92%, and 4.83% for CADEC, ShARe13, and ShARe14, respectively. However, despite not having the best overall scores in our experiments, the best DiscEnt models still outperform all of the baselines for the CADEC and ShARe14 datasets and are comparable to our overall best model highlighting the ability of our framework to extract discontinuous entities through word-pair triplets.
Table B1. Complete performance scores from the best-performing overall model for sentences with at least one discontinuous entity (DiscSent) and for discontinuous entities only (DiscEnt).
@@ Appendix B Comprehensive Metric Scores
Table B2. Complete performance scores from the best-performing discontinuous entity model for the overall dataset, for sentences with at least one discontinuous entity (DiscSent), and for discontinuous entities only (DiscEnt).
@@
Appendix C Token Gap Analysis
Figure C1 shows the difference in token gaps between CADEC, ShARe13, and ShARe14. CADEC generally shows shorter gaps between spans for discontinuous entities, while the ShARe datasets have wider gaps despite having shorter entities. These differences present unique challenges for extracting discontinuous entities in each dataset, highlighting the need for a flexible and adaptable solution like our proposed framework.
Figure C1. Distribution of token gaps of discontinuous entities.
Appendix D Hyperparameter Study
We investigate the effect of different hyperparameter values on our best overall model. In Table D1, we test different learning rate values for the Adam optimiser and find that the optimal learning rate value for our framework is 5e-04.
Table D1. Comparison of learning rates. Bold indicates best scores while underline shows next best.
@@ Appendix D Hyperparameter Study
Appendix E Best-found Parameter Setup
Table E1 outlines the best hyperparameter combinations for each dataset based on overall performance scores. PubMedBERT provides the best performance for both ShARe datasets while BioBERT works best for CADEC. A centroid triplet selection from the grid tag logits is optimal for CADEC and ShARe13 while a negative centroid selection from the word-pair grid is best for ShARe14. The window size varies for each dataset highlighting the importance of tuning these parameters.
Table E1. Parameter setup for the best model based on overall performance scores for each dataset.
@@ Appendix E Best-found Parameter Setup
Appendix F Large Language Model Prompts
Due to the increasing popularity of large language models (LLMs), we explored the application of out-of-the-box LLMs GPT-4o and Gemini for this task. We explored zero shot, one shot, few shot, and chain-of-thought (CoT) prompting techniques. Table F1 provides zero shot and few shot CoT templates while Table F2 provides the variables injected in the templates for each dataset. One Shot CoT prompt is similar to the Few Shot CoT except that only one example is provided. Non-CoT prompts remove the last line which asks the LLM to output an explanation.
Table F1. Prompt templates used for large language models.
@@ Appendix F Large Language Model Prompts
Zero Shot CoT ”The task is to find the index of the words from any {entity_descriptor} entities from the given text. The text input is already tokenized and is given in a list form where one entry corresponds to a word or punctuation. The word indexes must be based on the list. The entities may be continuous or discontinuous, single-word or multiple words. There may also be no entities in the text. Text: input Return the output in a json format followed by a set of steps to explain how the output was generated: “‘json [{\”entity\”: entity, \”index\”:[index1, index2 etc], \”type\”: \”{entity_type}\”}, {\”entity\”: entity, \”index\”:[index1, index2, index3 etc], \”type\”: \”{entity_type}\”}, etc]”’ Explanation: explanation”
Few Shot CoT ”The task is to find the index of the words from any {entity_descriptor} entities from the given text. The text input is already tokenized and is given in a list form where one entry corresponds to a word or punctuation. The word indexes must be based on the list. The entities may be continuous or discontinuous, single-word or multiple words. There may also be no entities in the text. Below are some examples of input text and output format. Input text: {input_example_1} Expected output: {output_example_1} Input text: {input_example_2} Expected output: {output_example_2} Now extract the entities from the text below following the examples above. Text: {input} Return the output in a json format followed by a set of steps to explain how the output was generated: “‘json [\”entity\”: entity, \”index\”:[index1, index2 etc], \”type\”: \”{entity_type}\”, \”entity\”: entity, \”index\”:[index1, index2, index3 etc], \”type\”: \”{entity_type}\”, etc]”’ Explanation: explanation”
Table F2. Variables and examples used for prompt engineering for each dataset.
@@
Appendix G Case Studies
We provide some case studies comparing our framework with the baseline model, Gemini, and GPT-4o. The green highlight shows a correct output, while the red highlight shows an incorrect output.
Figure G1. Case study for CADEC comparing results from trained models using our framework and a baseline and from zero and few-shot CoT prompt engineering using LLMs. The sample prompt provided follows the few-shot CoT template. All prompt templates are provided in Appendix F.
Figure G2. Case study for CADEC comparing results from trained models using our framework and a baseline and from zero and few-shot CoT prompt engineering using LLMs. The sample prompt provided follows the few-shot CoT template. All prompt templates are provided in Appendix F.
Figure G3. Case study for ShARe13 comparing results from trained models using our framework and a baseline and from zero and few-shot CoT prompt engineering using LLMs. The sample prompt provided follows the few-shot CoT template. All prompt templates are provided in Appendix F.
Figure G4. Case study for ShARe14 comparing results from trained models using our framework and a baseline and from zero and few-shot CoT prompt engineering using LLMs. The sample prompt provided follows the few-shot CoT template. All prompt templates are provided in Appendix F.